OpenStudy (anonymous):

how to differentiate : y = sqrt[ (1+tan x) / ( 1- tan x) ]

OpenStudy (loser66):

\[\sqrt{\dfrac{1+tanx}{1-tanx}}\] right?

OpenStudy (anonymous):

yes right

OpenStudy (loser66):

Let the sqrt aside, just calculate the terms inside, turn tan to sin/cos and simplify it what do you have?

OpenStudy (loser66):

tan = sin/cos

OpenStudy (anonymous):

\[\sqrt{\dfrac{\cos x + \sin x}{\cos x-\sin x}}\]